OpenStudy (anonymous):

a major league baseball team plays 160 games in a season. A team that wins about 60% of its games will usually make the playoffs. How many games does a team need to win to make the playoffs? h

OpenStudy (pfenn1):

# of games needed to win to make playoffs = (160 games/season) * (60 percent)

OpenStudy (pfenn1):

remember 60%=60/100=0.60

OpenStudy (anonymous):

can you please show the work? would appreciate it

OpenStudy (pfenn1):

# of games needed to make playoffs = (160) * (0.60) = 96 games